The Schock Manufacturing Company uses a standard cost system. The standards are set January 1 each year and remain unchanged until December 31 The standard costs set for the next year are: Diret materials S10.00 per unit Direct labor Overhead 7.50 per unit 600 per unit The labor standard above includes a methods change from 2 hours to 1% bours per unit, effective February 28, and a change in labor rates from $4.00 to $5,00 per hour effective February 28. Overhead will be applied on the basis of standard labor hours The standard overbead rate per hour is SA0 (1%, b unxsao-s600 per unit). The variatie por ion is200 per hour. The original budgeted production volume for January was 5,000 units The actual cost data for January are: 6.000 units $62,000 Units produced Direct materials used Direct labor Actual overhead incurred: Variable iead incurred: 100 hours 23,000 14.000 Fixed 1. Compute the variances for direct labor and variable overhead. 2. State the possible causes of each variance which you have computed. The following intormation applies to Wallace Shirt Cempany. Standard Data Per Unit Material-3 pieces at $2 per piece Labor-2 hours at $3 per hour Overbead: Fired-2 hours at $2.50 per hour (S50,000/10,000 units Variable-2 hours at $4 per hour 8 ($80,000/10.000 units) Actual Data Material-25,000 pieces of material were purchased at $210 per piece. Twenty-four thousand pieces of material were used. Labor-17,000 hours at $2.50 per hour Overhead-Variable, $69,000 fixed, $43,000 Actual units produced-9,000 Compute the following: (a) Materials purchase price variance (b) Materials quantity (usage) variance (c) Labor rate variance (4) Labor efficiency variance (e) Variable overbead spending variance (n Variable overhead eficiency variance
Answer #1

1.Direct Labor Cost Variance = Standard Cost – Actual Cost

= 8*6,000 – 46,000

= 2,000 F

Note: Standard Labor Cost per Unit for January = 2*4 = $8 per Unit

Direct Labor Rate Variance = (Standard Rate – Actual rate)*Actual Hours

= (4 – Actual Rate)*11,000

= 44,000 – 46,000

= 2,000 U

Direct Labor Efficiency Variance = (Standard Hours – Actual Hours)*Standard Rate

= (6,000*2 – 11,000)*4

= 4,000 F

Variable Overhead Cost Variance = Standard Cost – Actual Cost

=(4 *6,000 – 23,000)

= 1,000 F

Standard Rate = 2*2 = $4 per unit

Variable Overhead Rate Variance = (Standard Rate – Actual rate)*Actual Hours

= (2 – Actual Rate)*11,000

= 22,000 – 23,000

= 1,000 U

Variable Overhead Efficiency Variance = (Standard Hours – Actual Hours)*Standard Rate

= (6,000*2 – 11,000)*2

= 2,000 F

2.Reasons:

For Favorable Labor Cost/Spending Variance

1.Efficient Labor, fewer hours used than standard

For Favorable Variable Overhead Variance:

1.Efficient Labor, fewer hours used than standard

a)Material Purchase Price Variance = (Standard Rate – Actual Rate)*Actual Quantity

= (2-2.10)*25,000

= 2,500 U

b)Material Quantity/Usage Variance = (Standard Quantity – Actual Quantity)*Standard Price

= (9,000*3 – 24,000)*2

= 6,000 F

Note: Purchase Price Variance is calculated for all the units purchased while usage variance is calculated on the basis of units used

c)Labor rate Variance = (Standard Rate – Actual rate)*Actual Hours

=(3-2.5)*17,000

= 8,500 F

d)Labor Efficiency Variance = (Standard Hours – Actual Hours)*Standard Rate

= (9,000*2 – 17,000)*3

= 3,000 F

e)Variable Overhead Spending Variance = (Standard Rate - Actual Rate)*Actual Hours

= 4*17,000 – 69,000

= 1,000 U

f)Variable Overhead Efficiency Variance = (Standard Hours – Actual Hours)*Standard Rate

=(18,000 – 17,000)*4

= 4,000 F